Operations Report – March 22nd

Operations Report 22-MAR-2020

SOL: n/a

Name of person filing report: Shannon Rupert

Non-nominal systems: None

Notes on non-nominal systems: Nothing to report

Power system: SOLAR ONLY. Nominal.

Diesel Reading – 60%

Station Propane Reading – 68 %

Ethanol Free Gasoline – 5.5 gallons

Water (loft tank): full gallons

Water Meter: n/a units

Water (static tank) – 150 gallons

Static to Loft Pump used – no

Water in GreenHab – 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Perseverance rover used: Yes

Hours: 152.9

Beginning charge: 96 %

Ending charge: 100%

Currently charging: yes

Sojourner rover used: Yes

Hours: 105.2

Beginning charge: 100%

Ending charge: 55 %

Currently charging: yes

Spirit rover: Yes

Hours: 134.6

Beginning charge: 100%

Ending charge: 43 %

Currently charging: yes

Opportunity rover used: Yes

Hours: 70.7

Beginning charge: 100%

Ending charge: 47 %

Currently charging: yes

Curiosity rover used: Yes

Hours: 140.0

Beginning charge: 100%

Ending charge: 40 %

Currently charging: yes

Notes on rovers: All rovers were moved to town today for storage
ATV’s Used: No

Reason for use: n/a

Oil Added? no

ATV Fuel Used: None

Hours the ATVs were Used today: 0

Notes on ATVs: ATVs will be stored in a line outside the Hab until May.

HabCar used and why, where? Not used

CrewCar used and why, where? In Hanksville. Used today to travel
with the rovers and ferry people back to MDRS.

General notes and comments: Started loading some of the trash on the
trailer and hope to take another load to the dump this week.

Summary of internet: Nominal.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Worked on cleaning upper deck. Progress
made but I’m not finished

Summary of GreenHab operations: Still not quite done.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: The seal on the dome lip appears to
have shrunk. I sent photos to Peter and we are just going to keep an
eye on it for now and I will push it back into place as much as
possible when needed.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: Nothing at this time

Copyright © The Mars Society. All rights reserved. | Main Site